# MINDS: Detection of an inner gas disk caused by evaporating bodies around HD 172555

### v2

@@ 4.1.4 Summary of Gas Origin Scenarios
One important aspect requiring explanation is the stability of an inner gas disk at much smaller separations in HD 172555 than in Pic, despite the presence of FEBs in both systems. This favors a scenario in which gas is produced by material on close-in, low-eccentricity orbits, enabling the formation of a long-lived inner gas disk. Potential sources include a population of close-in bodies, the continuous outgassing of a large in-situ rocky planet, or the slow inward drift and sublimation of fine dust at the sublimation radius. For dust, the location of the sublimation radius strongly dependends on composition and particle size. FeS larger than 1.5 m would evaporate at radii of 0.1–0.3 au (see Appendix Fig. 7), comparable to the inferred gas disk location. Smaller particles would be vulnerable to blowout and would sublimate at larger separation (1 au). Johnson et al. (2012) suggested that 0.02 m-sized obsidian (SiO) grains would resist blowout, but these would not release the atomic species observed here. A detailed study remains needed to explore the combination of grain compositions, sizes, and morphologies, that would allow efficient inward drift, resistance to blowout, and sublimation at radii consistent with the observations.
The in-situ production scenario could resemble the recently observed catastrophically evaporating rocky planet BD+054868 Ab (Hon et al., 2025), or atmospheric loss from a sulfur-rich atmosphere, such as the super-Earth L98-59 d (Gressier et al., 2024; Banerjee et al., 2024). An in-situ rocky world could provide a long term supply of localized, high-density gas over long timescales, explaining the strong mid-IR atomic line emission and stable close-in gas disk.
Based on the observations presented here, we find that while transient events involving perturbed asteroid-like bodies from large radii could contribute episodically to gas production, they are unlikely to fully explain the stable inner gas disk observed in HD 172555. Similarly, in-situ collisional cascades are challenged by the short collisional timescales relative to the system age. However, this scenario cannot be fully ruled out, since the cascade could have been triggered recently in the lifetime of the system, potentially dynamically linked to the giant impact. Continuous gas production from a close-in evaporating rocky planet, and/or inward-drifting dust grains that evaporate at small radii remain plausible scenarios. Future observations and modeling of dust properties and gas dynamics will further constrain these scenarios.
@@ 5 Conclusions
This work presents the first detection of mid-infrared emission from an inner disk of hot atomic gas around HD 172555. The system is known for its warm dust belt in the terrestrial planet zone, thought to be the result of a giant impact. It is known to host star-grazing bodies (so-called falling evaporating bodies) seen via absorption lines and transit events. Our MIRI/MRS observations reveal mid-IR emission of nickel, chlorine, and sulfur for the first time in any debris disk, providing direct evidence that large amounts of rocky material are thermally disintegrating near the star and releasing iron and other elements into the gas phase.
We compared our discovery to HD 172555’s near-twin of the same moving group, Pic. Both systems share similar stellar properties yet differ in disk and planetary architecture, offering an excellent testbed for exploring how this gas relates to planet formation. The comparison shows that the inner gas disk in HD 172555 lies much closer to the star, despite the presence of FEBs in both systems. This suggests that the gas in HD 172555 is produced predominantly by an in-situ reservoir, rather than from material falling inward from larger separation.
We discuss four possible scenarios for the origin of the gas disk: (1) perturbed asteroid-like bodies originating from a closer-in reservoir; (2) an in-situ collisional cascades; (3) outgassing from a large body on a stable orbit, such as an evaporating rocky world; or (4) evaporation of dust drifting inward from the main debris belt. The presence of a close-in source of gas could plausibly be linked to the dynamical processes associated with the giant impact that is proposed to explain the warm dust belt. Such an event could have generated fragments and possibly led to enhanced dynamical activity in the inner system, including increased eccentricities and collisional interactions (Kral et al., 2015), or produced dust capable of drifting inward and sublimating (Su et al., 2020). However, these scenarios are challenging to distinguish and will require further multi-wavelength observations and more detailed modeling to fully disentangle.
@@
Our JWST data open up a valuable opportunity to directly constrain the chemical composition of rocky material in exoplanetary systems — an approach independent and complementary to measurements of polluted white dwarfs, and that can be compared to protoplanetary disk chemistry and evolution. Finally, we note that the observed gas and dust production in HD 172555, together with the detection of FEBs, supports the notion of ongoing dynamical stirring in the system, unless the observed gas is produced by an in-situ evaporating planet. Previous studies (e.g., Grady et al., 2018) have suggested that such stirring could be driven by the presence of planets. Our findings reinforce this picture, although a direct detection of planets in the system remains outstanding.
@@ 6 Acknowledgements
M.S., K.S. and T.H. acknowledge support from the European Research Council under the Horizon 2020 Framework Program via the ERC Advanced Grant Origins 83 24 28. B.T. acknowledges support from the Programme National PCMI of CNRS/INSU with INC/INP co-funded by CEA and CNES. M.T. acknowledge support from the ERC grant 101019751 MOLDISK. I.K., A.M.A., and E.v.D. acknowledge support from grant TOP-1 614.001.751 from the Dutch Research Council (NWO).
I.K., J.K., and T.K. acknowledge funding from H2020-MSCA-ITN-2019, grant no. 860470 (CHAMELEON). E.v.D. acknowledges support from the ERC grant 101019751 MOLDISK and the Danish National Research Foundation through the Center of Excellence “InterCat” (DNRF150). A.C.G. acknowledges support from PRIN-MUR 2022 20228JPA3A “The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 “NIR-dark Accretion Outbursts in Massive Young stellar objects (NAOMY)” and Large Grant INAF 2022 “YSOs Outflows, Disks and Accretion: towards a global framework for the evolution of planet forming systems (YODA)”. O.A. is a Senior Research Associate of the Fonds de la Recherche Scientifique – FNRS. OA and VC thank the European Space Agency (ESA) and the Belgian Federal Science Policy Office (BELSPO) for their support in the framework of the PRODEX Programme. V.C. acknowledges funding from the Belgian F.R.S.-FNRS. D.G. thanks the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA). G.P. gratefully acknowledges support from the Max Planck Society.
@@ Appendix A Detection of [S I] in Spitzer/IRS LH data
We have recalibrated the Spitzer/IRS LH (long wavelength, high-resolution) data taken on the 2004-03-22 (AOR 3563264) following the same approach as outlined in Sect. 2.1. For the high-resolution data the PSF-based extraction method is more reliable than aperture-based extraction, which was used for the low-resolution spectra. The sensitivity and spectral resolution is insufficient to detect any of the [Ni ii], [Cl I], and [Fe ii] lines. However, we have recovered the [S i] line at 25.249 micron, the strongest line in the spectrum. The line is detected in two orders (15 and 16) of the Spitzer/IRS LH data. Figure 4 shows the continuum subtracted emission line in JWST and the two IRS orders. The line is clearly detected in both orders. However, the width of the line is different. The line width indicates an instrument resolving power of R590 in order 15 and R350 in order 16. The expected resolving power according to the instrument documentation should be close to 600, which could indicate that the order 15 spectrum is more reliable. The integrated flux is measured in the same way as described in Appendix B.2 by fitting a linear trend (as the continuum) simultaneous to a (single) Gaussian. The window in which the fitting is done is 10 and 17 spectral channels wide for order 15 and 16, respectively, owing to different proximity to the edge of the order. The flux in order 15 is erg s-1 cm2. In order 16 it is erg s-1 cm2, driven by the larger width. The JWST measured flux is erg s-1 cm2. Overall the fluxes are similar between the Spitzer and JWST data, even though the observations are 20 years apart (2004-03-2 and 2023-09-18). If we assume that order 15 is more reliable, there may be a slight increase in the [S i] line flux in JWST compared to Spitzer, which could be related to the apparent slight brightening of the silicate feature between the Spitzer and JWST epochs.
@@ B.1 Continuum Subtraction
In order to reliably identify significant line emission in the spectrum we remove the stellar and dust continuum. We use the same method as Temmink et al. (2024). The default parameters were adjusted to include more spline nodes in the continuum, to reflect that we have an almost line-free spectrum with strong silica and silicate features. The quantile used to estimate the continuum was changed to 50%, as the spectrum scatters symmetrically around the continuum due to the lack of complex emission/absorption lines (contrary to protoplanetary disks).
@@ B.2 Line fitting and properties
The observed lines generally require two Gaussian components to fit and/or are broadened. We repeated the data reduction without Spec3 outlier rejection and residual fringe fitting steps (both on the detector and 1D spectrum) to make sure that the double peak is not a result of falsely rejecting the peak as an outlier. There was no impact on the shape of the features. However the amplitude did increase when masking out the strong emission peaks in the residual fringe correction step. We therefore used the masked reduction.
In order to measure the line fluxes, we simultaneously fit two Gaussians, plus an offset and slope to account for the local continuum, for all detected lines. The model used can be described by the following equation
where is the continuum flux at the reference wavelength , is the slope of the linear continuum, is the amplitude of each Gaussian emission line, is the standard deviation (width) of the Gaussians, is the offset from to each Gaussian center, and is the rest-frame wavelength of the emission line. The model is implemented using astropy-models and fitted with the gradient-descent optimization algorithm (LevMarLSQFitter). The fitted parameters are: , , , . The Gaussian standard deviation is fixed to the instrument resolution (except for [Fe II], which was better fit at a degraded ).
The fit was done within a window of 100 wavelength channels centered on the line. For determining the line flux, the amplitude of both Gaussian components are tied in the fit (), except for [S i] which has high S/N and shows signs of being asymmetric. The analytic integral is used to determine the line flux from the two Gaussian amplitudes and standard deviation
The uncertainty estimation is based on measuring the robust standard deviation (i.e the median absolute deviation multiplied by 1.4826, as implemented in astropy) of the residuals after subtracting the model fit in a 70 wavelength channel window, while excluding a window of 10 wavelength channels centered on the line. This residual scatter is used as the amplitude of Gaussians of the same line width to analytically estimate the uncertainty of the integrated flux (). This is valid under the assumption that the noise is homoscedastic, that is, all spectral channels have similar noise properties, and that the uncertainty of the width of the Gaussian is negligible compared to the uncertainty of the amplitude, thereby removing the cross-term as well.
For [Cl i] it is necessary to include the stellar HI 9–7 atomic hydrogen absorption line located at 11.309 m in the fit. Due to the fast stellar rotation it overlaps partially with the chlorine transition. In this case, we fit an additional broad negative Gaussian component.
@@ B.3 Resolving power
The instrument spectral resolving power of MIRI/MRS can change significantly within and between individual bands. For fitting lines and interpreting their properties, it is important to know for any given wavelength. We use the step-wise linear prescription for each MRS band derived in Pontoppidan et al. (2024) and refined for some of the bands in Banzatti et al. (2025). Where possible we have confirmed the resolving power derived from this prescription by measuring the line width of the same transition seen in HD 172555 in other MIRI/MRS data under the assumption that the lines are unresolved. All our transitions have been previously detected in atomic outflows of the protostars IRAS 23385+6053 (Gieser et al., 2023) and HH 211 (Caratti o Garatti et al., 2024). We measure all transitions in IRAS 23385+6053, except for [Cl I], which is not present. [Cl I] is measured in the spectrum of HH 211 instead. [Ni ii] at 6.6 m and [Fe ii] at 17.9 m appear to be broadened in both data sets, and we therefore assume the resolving power based Banzatti et al. (2025). The results are summarized in Table 3.
@@ B.4 Line widths and upper limits
Although the detected lines generally require two Gaussian components for a good fit, to measure the line width, we fit a single Gaussian component. The FWHM of all detected lines are given for HD 172555 in Table 4 and for Pic in Table 5.
@@
in Eq. B2. This gives an upper limit for non-zero velocity broadening of
@@ Appendix C Transition properties
Note. — is the Einstein coefficient, a measure of the transition probability. is the lower and the upper energy state of the transition. Transitions with are transitions to the ground state. in [K] is the energy of the upper level expressed in temperature units. Transitions above the horizontal line are detected in our HD 172555 data. Transitions below the horizontal line fall within the MIRI/MRS wavelength range, but are not significantly detected.
@@ Appendix D Keplerian line profiles
In order to confirm whether the observed broadened/double-peaked profiles can be explained with a rotating Keplerian gas disk, we have implemented a toy model similar to the one used by Schneiderman et al. (2021) to model the optically thin ALMA CO line observations for this target. In short, we draw a large number of samples from a uniform density distribution in a ring around the star. The ring has an inner and outer radius and an inclination. For the sampled points we compute the projected Keplerian velocity and compute a velocity histogram. We then transform the velocities into wavelength offsets with respect to the line center to get the expected line profile shape, which we convolve with the instrument resolution. Finally, we resample the model line profile to the MRS wavelengths to compare it to the data.
For all lines we assume the same disk inclination (), the best-fit from Schneiderman et al. (2021), which is also consistent with the scattered light observations of Engler et al. (2018), as this is a symmetric and optically thin model, which side of the disk is facing us does not matter. We also assume the same stellar mass M⊙. For Pic we use a stellar mass M⊙ and an inclination of .
We fit this line profile of the rotating disk model to the spectra using the MCMC package emcee (Foreman-Mackey et al., 2013) to derive posterior parameter distributions for the gas disk of each species.
Flat priors are used for all parameters, except for the resolving power , for which we use a Gaussian prior with a of 2%. A lower limit on equal to a stellar radius of or au (Kiefer et al., 2023) is imposed. is required to be larger than . The data uncertainties are scaled by a factor to match the mean value of the surrounding continuum scatter after continuum subtraction, unless this scaling factor is smaller than 1. In the fit, we allow the resolving power to vary by up to 5%, to account for remaining uncertainties in the derivation of the resolving power. Lastly, we fit an overall offset of the profile to account for potential systematic blue or red shifts.
@@
Appendix E Continuum subtracted spectrum and comparison to Pic
The continuum subtracted MIRI MRS spectra of HD 172555 and Pic are shown in Fig. 5, with a focus on wavelength regions containing interesting spectral features.

### v3

@@ 4.1.4 Summary of Gas Origin Scenarios
One important aspect requiring explanation is the stability of an inner gas disk at much smaller separations in HD 172555 than in Pic, despite the presence of FEBs in both systems. This favors a scenario in which gas is produced by material on close-in, low-eccentricity orbits, enabling the formation of a long-lived inner gas disk. Potential sources include a population of close-in bodies, the continuous outgassing of a large in situ rocky planet, or the slow inward drift and sublimation of fine dust at the sublimation radius. For dust, the location of the sublimation radius strongly depends on composition and particle size. FeS larger than 1.5 m would evaporate at radii of 0.1–0.3 au (see Appendix Fig. 7), comparable to the inferred gas disk location. Smaller particles would be vulnerable to blowout and would sublimate at larger separation (1 au). Johnson et al. (2012) suggested that 0.02 m-sized obsidian (SiO) grains would resist blowout, but these would not release the atomic species observed here. A detailed study remains needed to explore the combination of grain compositions, sizes, and morphologies, that would allow efficient inward drift, resistance to blowout, and sublimation at radii consistent with the observations.
The in situ production scenario could resemble the recently observed catastrophically evaporating rocky planet BD+054868 Ab (Hon et al., 2025), or atmospheric loss from a sulfur-rich atmosphere, such as the super-Earth L98-59 d (Gressier et al., 2024; Banerjee et al., 2024). An in situ rocky world could provide a long-term supply of localized, high-density gas over long timescales, explaining the strong mid-IR atomic line emission and stable close-in gas disk.
Based on the observations presented here, we find that while transient events involving perturbed asteroid-like bodies from large radii could contribute episodically to gas production, they are unlikely to fully explain the stable inner gas disk observed in HD 172555. Similarly, in situ collisional cascades are challenged by the short collisional timescales relative to the system age. However, this scenario cannot be fully ruled out, since the cascade could have been triggered recently in the lifetime of the system, potentially dynamically linked to the giant impact. Continuous gas production from a close-in evaporating rocky planet, and/or inward-drifting dust grains that evaporate at small radii remain plausible scenarios. Future observations and modeling of dust properties and gas dynamics will further constrain these scenarios.
@@ 5 Conclusions
This work presents the first detection of mid-IR emission from an inner disk of hot atomic gas around HD 172555. The system is known for its warm dust belt in the terrestrial planet zone, thought to be the result of a giant impact. It is known to host star-grazing bodies (so-called FEBs) seen via absorption lines and transit events. Our MIRI MRS observations reveal mid-IR emission of nickel, chlorine, and sulfur for the first time in any debris disk, providing direct evidence that large amounts of rocky material are thermally disintegrating near the star and releasing iron and other elements into the gas phase.
We compared our discovery to HD 172555’s near-twin of the same moving group, Pic. Both systems share similar stellar properties yet differ in disk and planetary architecture, offering an excellent testbed for exploring how this gas relates to planet formation. The comparison shows that the inner gas disk in HD 172555 lies much closer to the star, despite the presence of FEBs in both systems. This suggests that the gas in HD 172555 is produced predominantly by an in situ reservoir, rather than from material falling inward from larger separation.
We discuss four possible scenarios for the origin of the gas disk: (1) perturbed asteroid-like bodies originating from a closer-in reservoir; (2) an in situ collisional cascades; (3) outgassing from a large body on a stable orbit, such as an evaporating rocky world; or (4) evaporation of dust drifting inward from the main debris belt. The presence of a close-in source of gas could plausibly be linked to the dynamical processes associated with the giant impact that is proposed to explain the warm dust belt. Such an event could have generated fragments and possibly led to enhanced dynamical activity in the inner system, including increased eccentricities and collisional interactions (Kral et al., 2015), or produced dust capable of drifting inward and sublimating (Su et al., 2020). However, these scenarios are challenging to distinguish and will require further multiwavelength observations and more detailed modeling to fully disentangle.
@@
Our JWST data open up a valuable opportunity to directly constrain the chemical composition of rocky material in exoplanetary systems — an approach independent and complementary to measurements of polluted white dwarfs, and that can be compared to protoplanetary disk chemistry and evolution. Finally, we note that the observed gas and dust production in HD 172555, together with the detection of FEBs, supports the notion of ongoing dynamical stirring in the system, unless the observed gas is produced by an in situ evaporating planet. Previous studies (e.g., Grady et al., 2018) have suggested that such stirring could be driven by the presence of planets. Our findings reinforce this picture, although a direct detection of planets in the system remains outstanding.
@@ 6 Acknowledgements
M.S., K.S. and T.H. acknowledge support from the European Research Council under the Horizon 2020 Framework Program via the ERC Advanced Grant Origins 83 24 28. B.T. acknowledges support from the Programme National PCMI of CNRS/INSU with INC/INP co-funded by CEA and CNES. M.T. acknowledges support from the ERC grant 101019751 MOLDISK. I.K., A.M.A., and E.v.D. acknowledge support from grant TOP-1 614.001.751 from the Dutch Research Council (NWO).
I.K., J.K., and T.K. acknowledge funding from H2020-MSCA-ITN-2019, grant no. 860470 (CHAMELEON). E.v.D. acknowledges support from the ERC grant 101019751 MOLDISK and the Danish National Research Foundation through the Center of Excellence “InterCat” (DNRF150). A.C.G. acknowledges support from PRIN-MUR 2022 20228JPA3A “The path to star and planet formation in the JWST era (PATH)” funded by NextGeneration EU and by INAF-GoG 2022 “NIR-dark Accretion Outbursts in Massive Young stellar objects (NAOMY)” and Large Grant INAF 2022 “YSOs Outflows, Disks and Accretion: toward a global framework for the evolution of planet forming systems (YODA)”. O.A. is a Senior Research Associate of the Fonds de la Recherche Scientifique – FNRS. O.A. and V.C. thank the European Space Agency (ESA) and the Belgian Federal Science Policy Office (BELSPO) for their support in the framework of the PRODEX Programme. V.C. acknowledges funding from the Belgian F.R.S.-FNRS. D.G. thanks the Belgian Federal Science Policy Office (BELSPO) for the provision of financial support in the framework of the PRODEX Programme of the European Space Agency (ESA). G.P. gratefully acknowledges support from the Max Planck Society.
@@ Appendix A Detection of [S I] in Spitzer/IRS LH data
We have recalibrated the Spitzer/IRS LH (long wavelength, high-resolution) data taken on the 2004-03-22 (AOR 3563264) following the same approach as outlined in Sect. 2.1. For the high-resolution data, the PSF-based extraction method is more reliable than aperture-based extraction, which was used for the low-resolution spectra. The sensitivity and spectral resolution are insufficient to detect any of the [Ni ii], [Cl I], and [Fe ii] lines. However, we have recovered the [S i] line at 25.249 micron, the strongest line in the spectrum. The line is detected in two orders (15 and 16) of the Spitzer/IRS LH data. Figure 4 shows the continuum-subtracted emission line in JWST and the two IRS orders. The line is clearly detected in both orders. However, the width of the line is different. The line width indicates an instrument resolving power of R590 in order 15 and R350 in order 16. The expected resolving power according to the instrument documentation should be close to 600, which could indicate that the order 15 spectrum is more reliable. The integrated flux is measured in the same way as described in Appendix B.2 by fitting a linear trend (as the continuum) simultaneous to a (single) Gaussian. The window in which the fitting is done is 10 and 17 spectral channels wide for order 15 and 16, respectively, owing to different proximity to the edge of the order. The flux in order 15 is erg s-1 cm2. In order 16 it is erg s-1 cm2, driven by the larger width. The JWST measured flux is erg s-1 cm2. Overall, the fluxes are similar between the Spitzer and JWST data, even though the observations are 20 years apart (2004-03-2 and 2023-09-18). If we assume that order 15 is more reliable, there may be a slight increase in the [S i] line flux in JWST compared to Spitzer, which could be related to the apparent slight brightening of the silicate feature between the Spitzer and JWST epochs.
@@ B.1 Continuum Subtraction
In order to reliably identify significant line emission in the spectrum, we remove the stellar and dust continuum. We use the same method as Temmink et al. (2024). The default parameters were adjusted to include more spline nodes in the continuum, to reflect that we have an almost line-free spectrum with strong silica and silicate features. The quantile used to estimate the continuum was changed to 50%, as the spectrum scatters symmetrically around the continuum due to the lack of complex emission/absorption lines (contrary to protoplanetary disks).
@@ B.2 Line fitting and properties
The observed lines generally require two Gaussian components to fit and/or are broadened. We repeated the data reduction without Spec3 outlier rejection and residual fringe fitting steps (both on the detector and 1D spectrum) to make sure that the double peak is not a result of falsely rejecting the peak as an outlier. There was no impact on the shape of the features. However, the amplitude did increase when masking out the strong emission peaks in the residual fringe correction step. We therefore used the masked reduction.
In order to measure the line fluxes, we simultaneously fit two Gaussians, plus an offset and slope to account for the local continuum, for all detected lines. The model used can be described by the following equation:
where is the continuum flux at the reference wavelength , is the slope of the linear continuum, is the amplitude of each Gaussian emission line, is the standard deviation (width) of the Gaussians, is the offset from to each Gaussian center, and is the rest-frame wavelength of the emission line. The model is implemented using astropy-models and fitted with the gradient-descent optimization algorithm (LevMarLSQFitter). The fitted parameters are: , , , and . The Gaussian standard deviation is fixed to the instrument resolution (except for [Fe II], which was better fit at a degraded ).
The fit was done within a window of 100 wavelength channels centered on the line. For determining the line flux, the amplitudes of both Gaussian components are tied in the fit (), except for [S i], which has high S/N and shows signs of being asymmetric. The analytic integral is used to determine the line flux from the two Gaussian amplitudes and standard deviation
The uncertainty estimation is based on measuring the robust standard deviation (i.e., the median absolute deviation multiplied by 1.4826, as implemented in astropy) of the residuals after subtracting the model fit in a 70 wavelength channel window, while excluding a window of 10 wavelength channels centered on the line. This residual scatter is used as the amplitude of Gaussians of the same line width to analytically estimate the uncertainty of the integrated flux (). This is valid under the assumption that the noise is homoscedastic, that is, all spectral channels have similar noise properties, and that the uncertainty of the width of the Gaussian is negligible compared to the uncertainty of the amplitude, thereby removing the cross-term as well.
For [Cl i] it is necessary to include the stellar HI 9–7 atomic hydrogen absorption line located at 11.309 m in the fit. Due to the fast stellar rotation, it overlaps partially with the chlorine transition. In this case, we fit an additional broad negative Gaussian component.
@@ B.3 Resolving power
The instrument spectral resolving power of MIRI MRS can change significantly within and between individual bands. For fitting lines and interpreting their properties, it is important to know for any given wavelength. We use the step-wise linear prescription for each MRS band derived in Pontoppidan et al. (2024) and refined for some of the bands in Banzatti et al. (2025). Where possible, we have confirmed the resolving power derived from this prescription by measuring the line width of the same transition seen in HD 172555 in other MIRI MRS data under the assumption that the lines are unresolved. All our transitions have been previously detected in atomic outflows of the protostars IRAS 23385+6053 (Gieser et al., 2023) and HH 211 (Caratti o Garatti et al., 2024). We measure all transitions in IRAS 23385+6053, except for [Cl I], which is not present. [Cl I] is measured in the spectrum of HH 211 instead. [Ni ii] at 6.6 m and [Fe ii] at 17.9 m appear to be broadened in both data sets, and we therefore assume the resolving power based Banzatti et al. (2025). The results are summarized in Table 3.
@@ B.4 Line Widths and Upper Limits
Although the detected lines generally require two Gaussian components for a good fit, to measure the line width, we fit a single Gaussian component. The FWHMs of all detected lines are given for HD 172555 in Table 4 and for Pic in Table 5.
@@
in Eq. B2. This gives an upper limit for nonzero velocity broadening of
@@ Appendix C Transition properties
Note. — is the Einstein coefficient, a measure of the transition probability. is the lower and the upper energy state of the transition. Transitions with are transitions to the ground state. in [K] is the energy of the upper level expressed in temperature units. Transitions above the horizontal line are detected in our HD 172555 data. Transitions below the horizontal line fall within the MIRI MRS wavelength range, but are not significantly detected.
@@ Appendix D Keplerian line profiles
In order to confirm whether the observed broadened/double-peaked profiles can be explained with a rotating Keplerian gas disk, we have implemented a toy model similar to the one used by Schneiderman et al. (2021) to model the optically thin ALMA CO line observations for this target. In short, we draw a large number of samples from a uniform density distribution in a ring around the star. The ring has an inner and outer radius and an inclination. For the sampled points, we compute the projected Keplerian velocity and compute a velocity histogram. We then transform the velocities into wavelength offsets with respect to the line center to get the expected line profile shape, which we convolve with the instrument resolution. Finally, we resample the model line profile to the MRS wavelengths to compare it to the data.
For all lines, we assume the same disk inclination (), the best fit from Schneiderman et al. (2021), which is also consistent with the scattered light observations of Engler et al. (2018)—as this is a symmetric and optically thin model, which side of the disk is facing us does not matter. We also assume the same stellar mass M⊙. For Pic, we use a stellar mass M⊙ and an inclination of .
We fit this line profile of the rotating disk model to the spectra using the Markov Chain Monte Carlo package emcee (Foreman-Mackey et al., 2013) to derive posterior parameter distributions for the gas disk of each species.
Flat priors are used for all parameters, except for the resolving power , for which we use a Gaussian prior with a of 2%. A lower limit on equal to a stellar radius of or au (Kiefer et al., 2023) is imposed. is required to be larger than . The data uncertainties are scaled by a factor to match the mean value of the surrounding continuum scatter after continuum subtraction, unless this scaling factor is smaller than 1. In the fit, we allow the resolving power to vary by up to 5%, to account for remaining uncertainties in the derivation of the resolving power. Lastly, we fit an overall offset of the profile to account for potential systematic blue- or redshifts.
@@
Appendix E Continuum-subtracted spectrum and comparison to Pic
The continuum-subtracted MIRI MRS spectra of HD 172555 and Pic are shown in Fig. 5, with a focus on wavelength regions containing interesting spectral features.